(Please I dont know what to do I need help )
A prism has a volume of 100 cm3. If the length is tripled, the width is doubled, and the height remains the same, the new prism's volume will be _______ cm3.

Answer by ramkikk66(644)


Volume of a prism = Area of base * height
Here the base is a rectangle and area of base = length * width.
If the length is tripled and width is doubled, the new base area becomes
(3 * length)*(2 * width) = 6 * length * width = 6 * original area of base
Since height is unchanged, the new volume = new base area * height
= 6 * original base area * height
= 6 * original volume (===> since volume = base area * height)
= 6 * 100 = 600 cub-cm.
Hope you got it :)